OpenStudy (anonymous):

please check my answer? thanks!!!! ;) Solve. 3x2 - 10x + 5 = 0 x = 2 _+_ sq root 20

OpenStudy (anonymous):

\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]\[a=3,b=-10,c=5\]

OpenStudy (anonymous):

ugh:(

OpenStudy (anonymous):

plug and chug \[x=\frac{10\pm\sqrt{100-4\times 3\times 5}}{6}\]

OpenStudy (anonymous):

plug and chug hahaha ;)

OpenStudy (anonymous):

what else can you do? \[x=\frac{10\pm\sqrt{40}}{6}\] \[x=\frac{10\pm2\sqrt{10}}{6}\] \[x=\frac{5\pm\sqrt{10}}{3}\]

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

k

OpenStudy (anonymous):

x=2 sq root 10

OpenStudy (anonymous):

over 6

OpenStudy (anonymous):

dont forget the \(-b\) part in the numerator

OpenStudy (anonymous):

and don't forget the \(\pm\) part, there are two answers

OpenStudy (anonymous):

thank you ;)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

x = 5 + sq root 10/ 3? @satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

also \[\frac{5-\sqrt{10}}{3}\]